OpenStudy (anonymous):

A pair of equations is shown below: y = 6x − 4 y = 5x − 3 Part A: In your own words, explain how you can solve the pair of equations graphically. Write the slope and y-intercept for each equation that you will plot on the graph to solve the equations. (6 points) Part B: What is the solution to the pair of equations? (4 points)

OpenStudy (radar):

Part B: Note that both equations = y or y=y We can then write: 6x-4 = 5x -3 Now solve using algebra.

OpenStudy (radar):

Have you had algebra???

OpenStudy (anonymous):

pre algebra

OpenStudy (anonymous):

-4= 1x -3?

OpenStudy (er.mohd.amir):

6x-4=5x-3 6x-5x=4-3 solve

OpenStudy (anonymous):

0?

OpenStudy (er.mohd.amir):

x=1